Please find attached and below the first of many update reports for the mobilization of Ensign Rig 963 from Oates Rd, Houston to AMC, Perth.
We started the Quarantine operations on Monday the 2nd April.
As expected there were many teething problems but with the help and guidance of the Inspectors we managed to iron most of them out.
Monday and Tuesday were utilized to arrange the flow of cargo, stands, cleaning processes and overall plan of action.
Whilst the Pipetubs flew through the process, we ran into a solid wall once the first wrapped bundle of 40 ft matting was opened up on Tuesday.
There were insects, bugs, spiders, webs and mud on every piece. Due to the wooden content of the Matting, no Coarse cleaning could be conducted as they could not be inspected whilst wet.
We had to place stands at suitable positions in order to place one 40 ft matt on two stands to commence cleaning.
The week was mostly comprised of cleaning of 24 pieces of matting. These were the most time and space consuming.
In addition to the matting we have had 6 Std Pipetubs and 2 Hydraulic Pipetubs run the gauntlet and come out the other end with Red Tags affixed.
Currently only two Pipetubs have not been inspected, all other cargo that has entered the Quarantine Zone has been tagged and passed by the Inspectors.
24 pieces of 40ft Matting – Still awaiting Wood Treatment Certs.
4 Loader Accessible Pipetubs
2 Hydraulic Pipetubs
103 pieces of 3.5 inch DP
145 pieces of 5 inch DP
5 pieces of Loader Ramp
We have loaded suitable quantities of DP into Pipetubs to keep them within the legal weights for transport.
All matting that has been passed along with the Loader Ramp have been shrink wrapped.
We still need to shrink wrap the three Pipetubs that have been loaded with DP.
There are still 6 bundles of 2 pieces per bundle of 40 ft Matting along with 2 bundles of 3 pieces per bundle of 20 ft Matting that need to run the gauntlet.
Current working hours set in the port are Monday to Saturday 07h30 to 18h30. These times are subject to change.
We will be continuing with the 5 inch DP on Monday along with the shrink wrapping of the loaded Pipetubs.
Due to no cargo being available to be taken to the facility we had the inductions for all persons that will be involved on this project done on Monday the 2nd April at Boaz.
A brief run through of the process flow was then done.
It was made clear that whilst we had a plan of action, it was a working model and will be amended to suit the cargo as it arrives.
We started to see some cargo on the 3rd April with the two Site Storage Skids and Junk Skid being released to ITAC.
On the 4th and 5th we had a trailer load of loose items released and delivered to Boaz.
The two site storage containers and the 20ft container on the Junk skid have been Red Tagged, residual spray applied and sealed.
These are ready to be loaded with the loose items next week.
Cleaning, packing and combining of the loose items also commenced.
There is no work Sunday the 8th April.
We will commence again on Monday the 9th April at both facilities.